Find the cost required to carpet a hall 8m by 5.5 m if a rectangular section 2m by 3/2m is taken out to provide for the fire place ans the carpet costs $52.50 per m sq.
Answer by user_dude2008(1862)

2(3/2)=3
8*5.5 - 3 = 41

41*52.50 = 2152.5

total cost is $2152.50